4. Russo Cabinets is a custom cabinet builder. They recently completed a set of kitchen cabinets (Job Number 1887), as summarHI, I'm very lost on solving this problem since I do not know how to solve for the applied overhead cost when there is no written estimated overhead cost. Can someone demonstrate step through step of the problem?
Answer #1

It is given that russo applies Overhead costto jobs at a rate of $18 per direct labour hour

according to the sheet the total direct labour hours were 62 hours for job 1887

hence overhead cost = $18 per direct labour hour * 62 (direct labour hours)

= 18*62 =$1116 is the Overhead cost

Total Cost: Direct material cost+ labour cost+overhead

=1390+1482+1116 = $3988 is the toal cost